OpenStudy (anonymous):

Find the effective rate of interest for 6% compounded quarterly.

OpenStudy (amistre64):

would it be 6%/4 ??

OpenStudy (anonymous):

effective rate means what? simple interest rate with same interest?

OpenStudy (amistre64):

it means something, but i cant find a decent definition for it :)

OpenStudy (anonymous):

it is the equivalent annual simple rate of interest that would yield the same amount as compounding after one year

OpenStudy (amistre64):

Effective interest rate = (1 + r/n)^n -1 maybe?

OpenStudy (anonymous):

yes!!!! but do i need a principal amount? don't i?

OpenStudy (amistre64):

(1+.06/4)^4 = 1.06136355 .... 6.136...%

OpenStudy (amistre64):

the principal amount is just what tells you the amount of money you have; the interest rate doesnt really care about how much it is interest rating :)

OpenStudy (anonymous):

okay great!!! thank you so much!

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

so you just move the decimal over twice to get the answer?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

once you chop of the 1, yes; to convert a decimal format to percent format ... move the decimal over 2 spots

OpenStudy (anonymous):

\[(1+\frac{.06}{4})^4\] subtract 1, move the decimal twice to get the precent

OpenStudy (amistre64):

consider the decimal format as 1%; to get to 100% you multiply it by 100

OpenStudy (anonymous):

Find the effective rate of interest for 4% compounded quarterly. what about this one? the answer turns out to be 1^-8

OpenStudy (anonymous):

and i don't think that is correct....

OpenStudy (amistre64):

(1+.04/4)^4 -1 ... times 100

OpenStudy (anonymous):

ohhhh gotcha!!!!

OpenStudy (amistre64):

= 1.04060401 ; -1 and move decimal to get 4.0604... %